Required information Use the following information for the Exercises below. [The following information applies to the questions displayed below.] BMX Company has one employee. FICA Social Security taxes are 6.2% of the first $118,500 paid to its employee, and FICA Medicare taxes are 1.45% of gross pay. For BMX, its FUTA taxes are 0.6% and SUTA taxes are 2.9% of the first $7,000 paid to its employee. Gross Pay through August Gross Pay for September a. $ 6,800 $ 700 b. 18,600 2,500 c. 112,600 8,400 Exercise 9-5 Computing payroll taxes LO P2, P3 Compute BMX’s amounts for each of these four taxes as applied to the employee’s gross earnings for September under each of three separate situations (a), (b), and (c). (Round your answers to 2 decimal places.)
Answer #1
a)
Tax September Earnings Subject to Tax Tax Rate Tax Amount
FICA-Social Security 700 6.20% 43.40
FICA-Medicare 700 1.45% 10.15
FUTA 200 0.60% 1.20
SUTA 200 2.90% 5.80
b)
Tax September Earnings Subject to Tax Tax Rate Tax Amount
FICA-Social Security 2500 6.20% 155.00
FICA-Medicare 2500 1.45% 36.25
FUTA 0 0.60% 0.00
SUTA 0 2.90% 0.00
c)
Tax September Earnings Subject to Tax Tax Rate Tax Amount
FICA-Social Security 5900 6.20% 365.80
FICA-Medicare 8400 1.45% 121.80
FUTA 0 0.60% 0.00
SUTA 0 2.90% 0.00
Note: Earnings taxable for FUTA and SUTA in (a) = 7000-6800 = $200
Note: Earnings taxable for Social Security in (c) = 118500-112600 = $5900